OpenStudy (anonymous):

solve for x: 2/3 (x-2) = 4 - 2/5 - 5/2 2/5 5

OpenStudy (anonymous):

\[\frac{2}{3}(x - 2) = 4\] \[\frac{2}{3}x - \frac{4}{3} = 4\] \[\frac{2}{3}x = \frac{12}{3} + \frac{4}{3}\] \[\frac{2}{3}x = \frac{16}{3}\] \[\frac{3}{2} * \frac{2}{3}x = \frac{16}{3}*\frac{3}{2}\] \[x = 8\]

OpenStudy (anonymous):

not on the list of options. So i have a feeling i may of screwed something up -.-

OpenStudy (anonymous):

No it works, i dunno then.